OpenStudy (anonymous):

dividing a polynomial by a binomial. Please solve step by step. x2+8x+6/x+2

OpenStudy (amistre64):

divide it like you would normal numbers

OpenStudy (amistre64):

x +6 ------------ x+2| x^2 +8x +6 -x^2 -2x --------- 6x +6 -6x -12 --------- -6 <-- remainder

OpenStudy (anonymous):

\[x ^{2}+8x+6=x ^{2}+2x+6x+12-6=x(x+2)+6(x+2)-6\]=(x+2)(x+6)-6 \[x ^{2}+8x+6/(x+2)=((x+2)(x+6)-6)/(x+2)=(x+6) -6/(x+2)\]

OpenStudy (amistre64):

x +6 - [6/(x+2)]

OpenStudy (anonymous):

yes i mean that...